Find the mass of a silicon sample that's temperature increased by 29 oC when 382 J of heat were applied

sSi = .71 J/goC


1
Expert's answer
2021-03-10T06:53:40-0500

Q = C x m x deltaT

m = Q / (C x deltaT) = 382 / (0.71 x 29) = 18.55 g